1.) 4/5x=7/10





thanks-
Jordan

Found 2 solutions by checkley71, nycsharkman:
Answer by checkley71(8403)

4x/5=7/10
4x*10=5*7
40x=35
x=35/40
x=7/8 answer.

Answer by nycsharkman(136)

Solve for x
4/5x=7/10
The fraction 4/5x really is 4x/5 or (4/5)(x).
I will use 4x/5 = 7/10.
Cross-multiply.
4x times 10 = 5 times 7
40x = 35
Divide both sides of the equation by 40.
x = 35/40
Reduce fraction on right side to the lowest term.
x = 7/8